Solve using substitution:
-y − 5x = -7
2y − 2x = 14

Solve using substitution:
-8x = -4 − y
3y -24x = 8

Solve using elimination:
6x − 2y − 3= 0
4y − 4x = -6

Solve using elimination:
5x − 7y − 1= 0
-10y − 5x = -1

Solve using elimination:
-4y + 2x − 8 = 0
-3x + 6y = 12

Solve using elimination:
8y = -2x − 16
-x − 4y - 8 = 0


Answer by Alan3354(69443)

Solve using substitution:
-y − 5x = -7 --> y = -5x + 7
2y − 2x = 14 --> y - x = 7
Sub for y in the 2nd equation.
------------------------------
Solve using substitution:
-8x = -4 − y
3y -24x = 8
Do it like the 1st one.
---------------------------
Solve using elimination:
6x − 2y − 3= 0
4y − 4x = -6 --> -2x + 2y = -3
Add the 2 equaions to eliminate y terms.
-----------------------------
Solve using elimination:
5x − 7y − 1= 0
-10y − 5x = -1
Multiply the 1st eqn by 2, then add them.
---------------------------
Solve using elimination:
-4y + 2x − 8 = 0
-3x + 6y = 12
Solve using elimination:
8y = -2x − 16
-x − 4y - 8 = 0